# Single-step high-fidelity three-qubit gates by anisotropic chiral interactions

### v1

@@ V Conclusion
We introduced two new protocols for implementing a three-qubit controlled-controlled-rotation $\rm C^{2}Ry$ gate: (i) a single-step fully-synchronized protocol and (ii) a four-step protocol designed to further suppress systematic errors and noises.
@@ VI Acknowledgment
This research was supported by the EU through the H2024 QLSI2 project, by the Army Research Office under Award Number: W911NF-23-1-0110, and partly by the NCCR Spin Grant No. 51NF40-180604. The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Infeasibility of controlled-controlled-rotation gate with two-qubit interactions
We have established a no-go result for synchronizing general controlled-controlled-rotation gates. However, introducing a small transversal $ZZZ$ interaction resolves this issue, as explained in the main text. Given a synchronization condition $(m,n_{1},n_{2},n_{3})$ for Eq. (16), determining the synchronization ratios $J_{ij}/\Omega_{2}$ is best done numerically due to the complexity of the analytical expressions.

### v2

@@ V Conclusion
We introduced two new protocols for implementing a three-qubit controlled-controlled-rotation $\rm C^{2}Ry$ gate: (i) a single-step fully-synchronized protocol and (ii) a four-step protocol designed to further suppress systematic errors and noises. The protocols can be readily adapted to other three-qubit controlled-controlled-unitary gates.
@@ VI Acknowledgment
This research was supported by the EU through the H2024 QLSI2 project, by the Army Research Office under Award Number: W911NF-23-1-0110, and in part by NCCR Spin (grant number 225153). The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
@@ Appendix A Infeasibility of controlled-controlled-rotation gate with two-qubit interactions
We have established a no-go result for synchronizing general controlled-controlled-rotation gates. However, introducing a small longitudinal $ZZZ$ interaction resolves this issue, as explained in the main text. Given a synchronization condition $(m,n_{1},n_{2},n_{3})$ for Eq. (16), determining the synchronization ratios $J_{ij}/\Omega_{2}$ is best done numerically due to the complexity of the analytical expressions.
@@ Appendix B Microscopic model
The explicit interaction strength $J_{123}^{\parallel}$ in the ESOA frame is found to be proportional to
$J_{123}^{\parallel}U^{2}$ $\propto(t_{12,\uparrow}t_{23,\uparrow}t_{31,\uparrow}-t_{12,\downarrow}t_{23,\downarrow}t_{31,\downarrow})$ (63a)
$+(t_{13,\uparrow}t_{21,\uparrow}t_{32,\uparrow}-t_{13,\downarrow}t_{21,\downarrow}t_{32,\downarrow}),$ (63b)
which we identify with third-order tunneling events. The first (second) term describes the spin-dependent phase interference effect between spin-up and spin-down particles traveling counterclockwise (clockwise) to the triangular device in Fig. 1. By using Eq. (58), it is straightforward to find that the two terms are given by
$(t_{12,\uparrow}t_{23,\uparrow}t_{31,\uparrow}-t_{12,\downarrow}t_{23,\downarrow}t_{31,\downarrow})$ $=-2i\tau^{3}e^{i\Phi_{B}}\sin(3\bar{\gamma}),$ (64a)
$(t_{13,\uparrow}t_{21,\uparrow}t_{32,\uparrow}-t_{13,\downarrow}t_{21,\downarrow}t_{32,\downarrow})$ $=2i\tau^{3}e^{-i\Phi_{B}}\sin(3\bar{\gamma}).$ (64b)
The two terms differ by a minus sign due to time-reversal symmetry. Consequently, the combination of magnetic phase interference and spin-dependent phase interference give rise to the prefactor in the main text
$J_{123}^{\parallel}U^{2}\propto 4\tau^{3}\sin(\Phi_{B})\sin(3\bar{\gamma}).$ (65)